The Ferrell Transportation Company uses a responsi
Answer #1

The missing amounts for Planes:

Return on Investments = Controllable Margin / Average Operating assets.

Using the data provided, Controllable Margin i.e, C is 25,000,000 * 12% = 3,000,000. ( answer)

Contribution Margin = Controllable margin + controllable Fixed Cost

So, B is 3,000,000 + 1,500,000 = 4,500,000 ( answer)

Sales Revenue = Variable Costs + Controbution Margin

So, A is 5,500,000 + 4,500,000 = 10,000,000 ( answer)

-----------------------------------------

Solution for Taxis

(f) Average Operating assets = Conrollable Margin/ return on Investments = 80,000/10% = 800,000. ( answer)

(e) Controllable Fixed Cost = Contribution Margin - Controllable margin = 250,000- 80,000 = 170,000 ( answer)

(d) Variable Cost= Sales - Contribution Margin = 500,000 - 250,000 = 250,000.

_______________________________________

Solution for Limos

(i) Return on invest ments= Controllable Margin / Average operating assets = 210,000/1,500,000 = 14% ( answer)

(h) Controllable Fixed Cost = Contribution Margin- Controllable Margin = 480,000-210,000= 270,000 ( answer)

(g) Sales revenue = Variable cost + contribution margin = 300,000 + 480,000 = 780,000. (answer)